$SPAD/src/input images3.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- Color gallery pages 3-4
)r tknot
-- draw a seris of unfolding torus knots
for i in 0..4 repeat torusKnot(2, 2 + i/4, 0.5, 25, 250)

———-

References^{[1] nothing}

4